# Scalable multi-qubit intrinsic gates in quantum dot arrays

### v1

@@ II.2 Time evolution in the qubit frame
In general, the Hamiltonian governing dynamics of qubits can be split as $H=H_{0}+H_{\mathrm{I}}$, where $H_{0}$ is the intrinsic Hamiltonian, or self-energy, required for the proper definition of qubits, and $H_{\mathrm{I}}$ includes all other interactions and control signals. The intrinsic Hamiltonian is persistent and applies the unitary rotation $U_{0}(\tau)=\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0muH_{0}\tau}$ to the lab-frame quantum states. This unitary rotation defines the qubit frame, in which qubit states are stationary in absence of interaction, and in which all quantum gates are defined. The natural choice for our multi-dot array is $H_{0}=H_{\mathrm{dot}}$ and $H_{\mathrm{I}}=H_{\mathrm{ex}}$ as suggested by Eq. (5). Some literatures uses external microwave to define similar rotating frame. Although this approach is intuitive for a single-qubit system, it would be problematic if multiple qubits with different Zeeman energies are involved.
The qubit-frame time evolution $\widetilde{U}(\tau)$ is generated by the interaction picture Hamiltonian $U_{0}(\tau)^{\dagger}H_{\mathrm{I}}U_{0}(\tau)$, but the time-dependence makes direct integration difficult. Here, we take an equivalent approach by reversely rotating the qubit frame,
@@
We note that one cannot cancel the $H_{0}$ terms, which would require $H_{0}$ and $H_{\mathrm{I}}$ to commute. To proceed with the qubit-frame time evolution, we make the assumption that the system works in the non-degenerate regime, where the qubit ground and excited states are well-separated. Specially, the typical Zeeman energy should be much larger than the interdot exchange energy $\varepsilon^{\mathrm{Z}}\gg J$. Notably, this condition is also equivalent to $\lVert H_{0}\rVert\gg\lVert H_{\mathrm{I}}\rVert$, allowing the Hamiltonian $H=H_{0}+H_{\mathrm{I}}$ to be approached in a perturbative way.
[…]
$\widetilde{U}_{nn}(\tau)$ $=r_{nn}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E_{n}}+\sum_{m\neq n}r_{nm}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau(E^{\prime}_{mn}+\delta E_{n})}$ (8)
where $r_{nm}\equiv\lvert\langle n|\mathopen{}m^{\prime}\rangle\rvert^{2}$, $\delta E_{n}\equiv E^{\prime}_{n}-E_{n}$ and $E^{\prime}_{mn}\equiv E^{\prime}_{m}-E^{\prime}_{n}$. The derivation is exact up to this stage. Next we will make the perturbative assumption, which will allow us to split the sum in Eq. (8) into a major term with $r_{nn}\sim 1$, and many minor terms with $r_{nm}\sim O(J^{2})$. By keeping the leading-order effect, we make the approximation,
@@
$\widetilde{U}(\tau)\approx\operatorname{diag}(\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E^{(1)}_{1}},\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E^{(1)}_{2}},\cdots)\equiv\widetilde{U}_{\mathrm{ideal}},$ (9)
where $\delta E^{(1)}_{n}=\langle n|H_{\mathrm{ex}}|n\rangle$ are the first-order energy corrections. For the most part of this paper, we concern ourselves with the problems of what and how a useful quantum gate can be achieved with $\widetilde{U}_{\mathrm{ideal}}$. We attribute the difference between $\widetilde{U}(\tau)$ and $\widetilde{U}_{\mathrm{ideal}}$ as coherent errors, whose combined magnitude is should be of higher orders. We will examine these errors in details in Sec. III.3. In a sense, we can view quantum gates as first-order effects on the quantum dot array brought by exchange coupling. In comparison, the zeroth order effect defines the qubits while higher order effects are considered as errors.
@@ II.3 Grid and bond vectors
where $\bm{\Lambda}$ is the diagonal matrix defined by the first-order energy corrections, or simply the vector itself when its clear from the context. Using first-order perturbation theory, we can explicitly derive
@@
where we have introduced the shorthands $S_{w}$ and $T_{w}$ for the spin-conserved and spin-flipped tunneling strength.
Despite the simple form of each bond vector, it can be quite involved to calculate the Kronecker sum of all bonds in a large grid. In general, the vector components of $\bm{\Lambda}$ are summations of different $S_{w}$ and $T_{w}$ terms. The exact expression in terms of these bond strengths depends on the order of qubits and the geometry of the full grid, and there can have many possible ways of connection for a large collection of quantum dots. However, thanks to the structure of the each bond vector, the grid vector always satisfies the reflective symmetric condition,
@@ II.4 Free phase factors
Firstly, a global phase factor can be applied to any wave vector without inducing conceivable change of any physical observables. Hence two unitary maps related by an arbitrary global phase factor $\phi_{\mathrm{g}}$ are completely equivalent, $\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi_{\mathrm{g}}}U\doteq U$, where the dot-equal symbol is used to represent such equivalence. More importantly, we also allow two unitary gates to differ by local phase gates on individual qubits. The rationale behind such local phase freedom is that $Z$-axis rotations can be implemented by the so-called virtual-$Z$ gates in the software level and do not require actual operations on the qubits [61, 62]. This local phase gate freedom is also exploited in original the proposal of two-qubit CZ gate [23]. In this paper, we will assume that such single-qubit $Z$ rotations come at trivial cost of state fidelity or operation time. The phase gate for a single qubit is the operator $Z_{\phi}=\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu(\phi/2)\sigma_{z}}\doteq\operatorname{diag}(1,\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi})$. Direct product of single qubit phase gates gives local phase gates for multi-qubits. Combing the global and local phase degree of freedom, we can define a free phase map
@@
where the gate vector $\bm{\Theta}_{\mathrm{G}}$ consists of real numbers associated with the gate. Focusing only on the phase factors, we obtain from Eqns. (9, 14, 17) the relation for the vectors,
@@ III Intrinsic Multi-qubit gates
Equipped with the Hamiltonian and phase corrections, in this section we study what multi-qubit gates can be intrinsically achieved on a general spin-qubit grid. We will first consider some specific examples and their potential applications then gives a general theory for identifying the accessible quantum gates.
@@ III.1 Multi-qubit controlled-phase gates
A particular useful class of quantum gates is the general controlled-phase gates. We require these gates to involve at least one control qubit and at least one target qubit subjecting to phase-shift according to the state of the control qubit(s). Notable examples include the regular controlled-Z (CZ) gate, the Toffoli-type controlled-controlled-Z (CCZ) gate and the parity checker controlled-Z-Z (CZZ) gate. Here we will examine of what gates are achievable and in what kind of geometry can it be achieved.
@@ III.1.1 Composition rules
Without loss of generality, we fix the first qubit as the control qubit and label it by ‘$\mathsf{C}$’, with the rest qubits, which may or may not be involved in the gate, labelled by $11,2,\cdots,n\equiv N_{Q}-$ 1. The full Hilbert space is thus spanned by $\{|0\rangle,|1\rangle\}_{\mathsf{c}}\otimes\{|0\rangle,|1\rangle\}_{1}\otimes\cdots\otimes\{|0\rangle,|1\rangle\}_{n}$. Setting the first qubit as control implies that the first half of the computational states remain intact under the gate transformation, while the second half is subject to certain phase flips. The gate vector of such controlled-phase gate can be written as
@@ III.1.3 General multi-qubit control gates
Let us now examine the the dynamics rule for the multi-qubit control gate. As hinted by the three qubit example, it turns out that such gate can be generally implemented in the stellar geometry. In the stellar geometry, the control qubit lives in the center of a cluster, and all other qubits are directly connected to the control qubit via bonds $1,2,\cdots,n$. The reduced grid vector in such case can be worked out as
@@
This solution suggests that such multi-qubit control is also allowed dynamically. This completes the full proof that the multi-qubit controlled phase gates are intrinsic. Of course, we have hitherto only considered the theoretical perspectives and leave out the technical details. From Eq. (36), we see that arbitrary controlled phases can be in principle achieved by controlling the tunnelling strength $T_{j}$ and $S_{j}$ for each bond. But as the qubit number scales up, the resources for simultaneously calibrating and controlling all bond can quickly become unmanageable. We will discuss a more robust and scalable scheme to fulfill this dynamics condition in Section IV.
We can also examine the multi-qubit controlled-phase under another type of geometry—the linear geometry. In the linear geometry, the qubits are labeled sequentially from one end to another end of an 1D chain (in a topologically equivalent way), and are connected with near-neighbors. The linear geometry cannot implement a general multi-qubit controlled-phase gate. But there is one exception: For a homogeneous chain with one control and $n$ target qubits, a $\pi$-phase can accumulate on both ends $\phi_{\mathsf{c}}=\phi_{n}=\pi$ but coherently cancels out for other qubits $\phi_{1}=\phi_{2}=\cdots\phi_{n-1}=0$. Such solution is already discussed in the three-qubit example. We can prove the general case with mathematical induction by noting the following recurrence relation
@@
where $\bm{\lambda}^{(n)}$ is the the grid vector for chains with $n$ target qubits. In general, the solution to the dynamics rule can be found for
@@
whence the resulting uncorrected time evolution is simply
@@
We remark that the superexchange oscillations observed for boundary states of a spin chain are manifestation of this gate [42, 43]. Such gate might be trivial from the quantum computational perspective, as it is just joint single-qubit $Z$. But the underlying physics is quite interesting. The final gate does not depend on the length and participating qubits involved in the path connecting the control and the final target qubit. In other words, it has the property of being topologically invariant.
@@ III.2 General gate decompositions
The MQCP gates stands for only a small fraction of what can be intrinsically achieved on a multi-qubit array. And there can be many different intrinsic gates under different types of dot connectivity. But in fact, all intrinsic gates can be decomposed into simultaneous MQCP gates. As a result, all the conclusions regarding local phase corrections, gate fidelity, and scalable implementations can be directly carried over.
@@
Proper local phase corrections are still required for a multi-qubit array to implement a product of MQCP gates. It is straightforward to show that the necessary phase corrections are summations of the phase corrections for each MQCP subgroups. This also explains the local phase identity $\phi_{\mathsf{c}}=\sum_{j}\phi_{j}$ for a general MQCP gate, which can be decomposed into multiple two-qubit groups with $\phi_{\mathsf{c}}=\phi_{j}$. While for more general arrays, there can have multiple viable ways of gate decomposition. Consider the example in Fig. 2(c), we have two different ways of decomposition into MQCP gates. At first sight, this ambiguity seems to give rise to uncertainties in local phase corrections. But it turns out the local phase corrections for a particular qubit only depends on the total tunneling strengths of the bonds connecting to that qubit. Therefore, different ways of decompositions are just equivalent ways of understanding the same gate. For the Fig. 2(c) example, assuming that all bonds are homogeneous and satisfy the condition $(T-S)\tau=\pi/2$, we find $\phi_{1}=\phi_{2}=3\pi/2$ and $\phi_{3}=\phi_{4}=\phi_{5}=\phi_{6}=\pi/2$ for both decompositions. In general, the necessary local phase factors are found by
$\phi_{j}=\sum_{j\in w}\tau(T_{w}-S_{w}).$ (41)
III.3 Fidelity estimations
We have hitherto replaced the actual time evolution in the rotating frame with the ideal time evolution $\widetilde{U}(\tau)\approx\widetilde{U}_{\mathrm{ideal}}(\tau)$ and considered what can be achievable with $\widetilde{U}_{\mathrm{ideal}}$. The difference between the actual and ideal unitarity maps is regarded as coherent error. In this section, we examine how much error is brought by making such approximation.
In general, we can characterize quantum gate implementations using the average gate fidelity, defined by averaging the fidelity for output states over all input states. Here, we are concerned with estimating the fidelity loss by replacing $\widetilde{U}$ with $\widetilde{U}_{\mathrm{ideal}}$. As both are unitary maps with the same dimensionality, the fidelity function can be calculated by

### v2

@@ II.2 Time evolution in the qubit frame
For any quantum computing system, the Hamiltonian governing qubit dynamics can be split as $H=H_{0}+H_{\mathrm{I}}$, where $H_{0}$ is the intrinsic Hamiltonian (self-energy) for proper definition of qubits; $H_{\mathrm{I}}$ includes the external control and internal coupling necessary for manipulation of the register states. $H_{0}$ is typically static and induces the unitary transform $U_{0}(\tau)=\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0muH_{0}\tau}$ to the lab-frame states. This constant rotation defines the qubit frame and qubit states. A quantum gate is a linear map among qubit states. It is implemented by the qubit-frame time evolution $\widetilde{U}(\tau)$ generated by $\widetilde{H}_{\mathrm{I}}(\tau)=U_{0}(\tau)^{\dagger}H_{\mathrm{I}}U_{0}(\tau)$. Usually, the time-dependence in $\widetilde{H}_{\mathrm{I}}(\tau)$ makes time integration difficult to solve analytically. We can take an equivalent approach by reversely rotating the qubit frame and express $\widetilde{U}(\tau)$ as the product
@@
We note that one cannot cancel the $H_{0}$ terms, which would require $H_{0}$ and $H_{\mathrm{I}}$ to commute. For the spin-qubit chip hosted on a multi-dot array, we set $H_{0}=H_{\mathrm{qu}}$ and $H_{\mathrm{I}}=H_{\mathrm{ex}}$, as naturally suggested by Eq. (6). It is also possible to use the external microwave frequency to define similar rotating frame, but such choice would be problematic if multiple qubits with different Zeeman energies are involved [62].
@@
To proceed with the qubit-frame time evolution, we make the assumption that the typical Zeeman energy is much larger than the typical interdot exchange energy $\varepsilon^{\mathrm{Z}}\gg J$ in the device working regime. Notably, this condition suggests $\lVert H_{\mathrm{qu}}\rVert\gg\lVert H_{\mathrm{ex}}\rVert$ and allows the Hamiltonian $H=H_{\mathrm{qu}}+H_{\mathrm{ex}}$ to be approached in a perturbative way. It is obvious form Eq. (6) that the $H_{\mathrm{qu}}$ eigenstates are just $\{|n\rangle\}$, with corresponding eigenenergies
$E_{n}=\sum_{n}\frac{1}{2}\operatorname{sig}(\sigma_{n_{j}})\varepsilon^{\mathrm{Z}}_{j}.$ (9)
Let us denote the eigenstates and eigenenergies of $H$ as $\{|n^{\prime}\rangle\}$ and $\{E_{n}^{\prime}\}$, which are perturbed from the eigenstates $\{|n\rangle\}$ and eigenenergies $\{E_{n}\}$ of $H_{\mathrm{qu}}$. The matrix exponential of $H_{0}$ and $H$ for $\widetilde{U}(\tau)$ can be taken under the respective eigen-basis $\{|n\rangle\}$ and $\{|n^{\prime}\rangle\}$. In particular, the diagonal elements of $\widetilde{U}(\tau)$ can be found by
$\langle n|\widetilde{U}(\tau)|n\rangle$ $=r_{nn}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E_{n}}+\!\sum_{m\neq n}r_{nm}\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau(E^{\prime}_{m}-E_{n})},$ (10)
where $r_{nm}\equiv\lvert\langle n|\mathopen{}m^{\prime}\rangle\rvert^{2}$ and $\delta E_{n}\equiv E^{\prime}_{n}-E_{n}$. While the expression for Eq. (10) is exact, the perturbative assumption suggests that it consists of a major term with $r_{nn}\sim 1-O(J^{2})$, and a summation of minor terms with $r_{nm}\sim O(J^{2})$, for $m\neq n$. As a result, we construct the following “ideal” gate $\widetilde{U}_{\mathrm{ideal}}\approx\widetilde{U}$, where only the first-order effects are kept,
@@
$\widetilde{U}_{\mathrm{ideal}}\equiv\operatorname{diag}(\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E^{(1)}_{1}},\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E^{(1)}_{2}},\cdots,\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\tau\delta E^{(1)}_{d}}),$ (11)
where $\delta E^{(1)}_{n}=\langle n|H_{\mathrm{ex}}|n\rangle\sim O(J)$ is the first order energy correction for the $n$th level. In this sense, we can view the intrinsic gates on quantum dot arrays as the first-order dynamical effects of the exchange coupling. In comparison, the zeroth-order term defines the qubit Hilbert space, while all the higher-order terms are attributed as errors.
We should remark that the identity $\delta E^{(1)}_{n}=\langle n|H_{\mathrm{ex}}|n\rangle$ is a result of non-degenerate perturbation theory, where the $E_{n}$ values are assumed to be distinctive. Thus it is tempting to also impose the requirements $\lvert E_{n}-E_{m}\rvert>\lVert H_{\mathrm{ex}}\rVert$ for all $n,m$. However, such constraints are neither practical for large-scale systems nor necessary for the definition of $\widetilde{U}_{\mathrm{ideal}}$. The difference between $\widetilde{U}(\tau)$ and $\widetilde{U}_{\mathrm{ideal}}$ is viewed as a source of coherent error in gate implementation. In the case of near-degenerate $E_{n}$ values, the perturbative series converges slowly and approximating $\widetilde{U}(\tau)$ by $\widetilde{U}_{\mathrm{ideal}}$ will necessarily result in large error. We will estimate the coherent error and the suppression of it in further details in Sec. III.3. For the most part of this paper, we focus on what and how a useful quantum gate can be achieved with $\widetilde{U}_{\mathrm{ideal}}$.
@@ II.3 Grid and bond vectors
where $\bm{\Lambda}$ is the diagonal matrix defined by the first-order energy corrections, or simply a vector of diagonal elements when its clear from the context. Using first-order perturbation theory, we can explicitly derive
@@
where we have introduced the shorthands $S_{w}$ and $T_{w}$ for the spin-conserved and spin-flipped tunneling strengths.
Despite the simple form of each bond vector, it can be quite involved to calculate the Kronecker sum of all bonds in a large grid. The vector components of $\bm{\Lambda}$ are summations of different $S_{w}$ and $T_{w}$ terms. Their exact expressions depend on the order of qubits and the connectivity of the full grid, and there can have many possible ways of connection for a large collection of quantum dots. However, thanks to the structure of the each bond vector, one can show that the grid vector always satisfies the reflective symmetric condition
@@ II.4 Free phase factors
Firstly, a global phase factor can be applied to any wave vector without inducing conceivable change of any physical observables. Hence two unitary maps related by an arbitrary global phase factor $\phi_{\mathrm{g}}$ are completely equivalent, $\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi_{\mathrm{g}}}U\doteq U$, where the dot-equal symbol is used to represent such equivalence. More importantly, we also allow two unitary gates to differ by local phase gates on individual qubits. The rationale behind such local phase freedom is that $Z$-axis rotations can be implemented by the so-called virtual-$Z$ gates in the software level and do not require actual operations on the qubits [62, 63]. This local phase freedom is also exploited in original the proposal of two-qubit CZ gate [23]. In this paper, we will assume that such single-qubit $Z$ rotations come at trivial cost of state fidelity or operation time. The phase gate for a single qubit is the operator $Z_{\phi}=\mathrm{e}^{-\mkern 1.0mu\mathrm{i}\mkern 1.0mu(\phi/2)\sigma_{z}}\doteq\operatorname{diag}(1,\mathrm{e}^{\mkern 1.0mu\mathrm{i}\mkern 1.0mu\phi})$. Direct product of single qubit phase gates gives local phase gates for multi-qubits. Combing the global and local phase degree of freedom, we can define a free phase map
@@
where the gate vector $\bm{\Theta}_{\mathrm{G}}$ consists of real numbers associated with the gate. Focusing only on the phase factors, we obtain from Eqs. (11, 16, 19) the relation for the vectors,
@@ III Intrinsic Multi-qubit gates
Equipped with the Hamiltonian and phase corrections, in this section we study what multi-qubit gates can be intrinsically achieved on a general spin-qubit grid. We will first consider some specific examples and their potential applications then develop a general theory for identifying the intrinsic quantum gates.
@@ III.1 Multi-qubit controlled-phase gates
A particular useful class of quantum gates is the general controlled-phase gates. We require these gates to involve at least one control qubit and at least one target qubit subjecting to phase-shift according to the state of the control qubit(s). Notable examples include the regular controlled-Z (CZ), the Toffoli-type controlled-controlled-Z (CCZ) and the parity checker controlled-Z-Z (CZZ) gate. Here we will examine of what gates are achievable and in what kind of geometry can it be achieved.
@@ III.1.1 Composition rules
Without loss of generality, we fix the first qubit as the control qubit and label it by ‘$\mathsf{C}$’, with the rest qubits, which may or may not be involved in the gate, labelled by $11,2,\cdots,n\equiv N_{Q}-$ 1. Setting the first qubit as control implies that the first half of the computational states remain intact under the gate transformation, while the second half is subject to certain phase flips. Therefore the gate vector of such gate can be written as
@@ III.1.3 General multi-qubit control gates
Let us now examine the the dynamics rule for the multi-qubit control gate. As hinted by the three qubit example, it turns out that such gate can be generally implemented in a stellar geometry. In the stellar geometry, the control qubit lives in the center of a cluster, and all other qubits are directly connected to the control qubit via bonds $1,2,\cdots,n$. The reduced grid vector in such case can be worked out as
@@
This solution suggests that such multi-qubit control is also allowed dynamically. This completes the full proof that the multi-qubit controlled phase gates are intrinsic. Of course, we have hitherto only considered the theoretical perspectives and leave out the technical details. From Eq. (38), we see that arbitrary controlled phases can be in principle achieved by controlling the tunnelling strength $T_{j}$ and $S_{j}$ for each bond. But as the qubit number scales up, the resources for simultaneously calibrating and controlling all bonds quickly become unmanageable. We will discuss a more robust and scalable scheme to fulfill this dynamics condition in Section IV.
We can also examine the multi-qubit controlled-phase under another type of geometry—the linear geometry. In the linear geometry, the qubits are labeled sequentially from one end to another end of an 1D chain (in a topologically equivalent way), and are connected with near-neighbors. A general multi-qubit controlled-phase gate cannot be implemented in the linear geometry. But there is one exception: For a homogeneous chain with one control and $n$ target qubits, a $\pi$-phase can accumulate on both ends $\phi_{\mathsf{c}}=\phi_{n}=\pi$ but coherently cancels out for other qubits $\phi_{1}=\phi_{2}=\cdots\phi_{n-1}=0$. Such solution is already discussed in the three-qubit example. We can prove the general case with mathematical induction by noting the following recurrence relation
@@
where $\bm{\lambda}^{(n)}$ is the the grid vector for chains with $n$ target qubits. In such case, the solutions to the dynamics rule exist at
@@
The resulting uncorrected time evolution is simply
@@
We remark that the superexchange oscillations observed for boundary states of a spin chain are manifestation of this gate [42, 43]. Such gate might be trivial from the quantum computational perspective, as it is just joint single-qubit $Z$. But the underlying physics is quite interesting. The final gate does not depend on the shape and number of qubits involved in the path connecting the control and the final target qubit. In other words, it has the property of being topologically invariant.
@@ III.2 General gate decompositions
The MQCP gates stand for only a small fraction of what can be intrinsically achieved on a multi-qubit array. And there can be many different intrinsic gates under different types of dot connectivity. But it turns out that all the intrinsic gates can be decomposed into simultaneous MQCP gates. As a result, all the conclusions regarding local phase corrections, gate fidelity, and scalable implementations can be directly carried over.
@@
Proper local phase corrections are still required for a multi-qubit array to implement a product of MQCP gates. It is straightforward to show that the necessary phase corrections of a dot should be accumulate from each MQCP subgroups. This also explains the local phase identity $\phi_{\mathsf{c}}=\sum_{j}\phi_{j}$ for a general MQCP gate, which can be decomposed into multiple two-qubit groups with $\phi_{\mathsf{c}}=\phi_{j}$. While for more general arrays, there can have multiple viable ways of gate decomposition. Consider the example in Fig. 2(c), we have two different ways of decomposition into MQCP gates. At first sight, this ambiguity seems to give rise to uncertainties in local phase corrections. But it turns different ways of decompositions are just equivalent ways of understanding the same gate. For the Fig. 2(c) example, assuming that all bonds are homogeneous and satisfy the condition $(T-S)\tau=\pi/2$, we will have $\phi_{1}=\phi_{2}=3\pi/2$ and $\phi_{3}=\phi_{4}=\phi_{5}=\phi_{6}=\pi/2$ for both decompositions. In general, the necessary local phase factors are found by
$\phi_{j}=\sum_{j\in w}\tau(T_{w}-S_{w}),$ (43)
that is, the necessary local phase correction for a dot is the summation of the accumulated phase of its connecting edges. Therefore, the local phase corrections are invariant to different ways of MQCP decomposition.
III.3 Estimation of the coherent error
We have hitherto replaced the actual rotating-frame map $\widetilde{U}(\tau)$ with the “ideal” map $\widetilde{U}_{\mathrm{ideal}}$ and investigated what can be achievable with the latter. However, the conditions for $\widetilde{U}_{\mathrm{ideal}}=G$ for a certain gate $G$ do not imply that the same gate can be accurately implemented with $\widetilde{U}(\tau)$. Difference between the actual and ideal maps is a coherent error in the gate implementation. In this section, we examine how much error is brought by making such approximation.
III.3.1 Fidelity bound
A implementation of quantum gate can be characterized with the average gate fidelity, defined by averaging the fidelity for output states over all input states. Here, we are concerned with estimating the fidelity loss by replacing $\widetilde{U}$ with $\widetilde{U}_{\mathrm{ideal}}$, both are unitary maps with the same dimensionality. The fidelity can be calculated by